OpenStudy (anonymous):

A boy drops a coin off a 64-foot cliff. If the equation of the height, h, of the coin t seconds after it drops is h = -16t2 + 64, about how many seconds will the coin take to reach the bottom of the cliff? 1 second 1.5 seconds 2 seconds 2.5 seconds

OpenStudy (anonymous):

To be frank I dont think the equation is quite right.

OpenStudy (anonymous):

lol that wouldnt be a first.... book is riddled with errors.

OpenStudy (anonymous):

Why am I not surprised :D

OpenStudy (anonymous):

Any ida how this is suppose to go?

OpenStudy (anonymous):

If you substitute h=64, you get t=0. Thats baseless coz the coin would be on the ground the whole time. Well try other problems.

OpenStudy (anonymous):

Solve the equation by graphing. x2 - 18x + 81 = 0 {-2, 9} {9/2, 3} {9} no real solution What about this one?

OpenStudy (anonymous):

9 is the ans

OpenStudy (anonymous):

Solve the equation by graphing. x2 + 5x + 8 = 0 {-5, 8} {2, 4} {-4, 5} no real solution

OpenStudy (anonymous):

Nope no solution

OpenStudy (anonymous):

Solve the equation by graphing. 3x2 + 9x = 12 {-4, 1} {-6, 2} {-3, 4} no real solution

OpenStudy (anonymous):

{-4,1}

OpenStudy (anonymous):

Solve the equation by graphing. x2 - 5x - 24 = 0 {4, 6} {-2, 12} {-4, 6} {-3, 8}

OpenStudy (anonymous):

{-3,8}

OpenStudy (anonymous):

Estimate the roots of the equation by graphing. Round to the nearest tenth and write in set form, such as {-1.2,3.4}. x2 - 6= 0